5^-2=
a) 1/25
b)-1/25
c)1/10
d)-10

Answer by stanbon(75887)

5^-2=
The negative in the exponent mens "invert".
The 2 in the exponent means "square".
You may do those operations in any order, but you have to perform both actions.
----------------
If you "invert" 1st you get (1/5)^2 = 1/25
If you "square" 1st you get (25)^-1 = 1/25
-----------------
----
a) 1/25
b)-1/25
c)1/10
d)-10
----------------
cheers,
Stan H.